A group of 2175 students were surveyed about the
courses they were taking at their college with
the following results:
1025 students said they were taking Dance.
1087 students said they were taking History.
1115 students said they were taking Math.
585 students said they were taking History and Math.
501 students said they were taking Math and Dance.
537 students said they were taking History and Dance.
243 students said they were taking all three courses.

Found 5 solutions by mccravyedwin, ikleyn, Edwin McCravy, greenestamps, MathTherapy:
Answer by mccravyedwin(407)

A group of 2175 students were surveyed about the courses
they were taking at their college with the following results:
Put the 8 clues in reverse order, for that's the order you need them.

1.  243 students said they were taking all three courses.
2.  537 students said they were taking History and Dance.
3.  501 students said they were taking Math and Dance.
4.  585 students said they were taking History and Math.
5.  1115 students said they were taking Math.
6.  1087 students said they were taking History.
7.  1025 students said they were taking Dance.
8.  2175 students were surveyed 

Draw a Venn diagram with three overlapping circles, labeled D, H, and M, 
all inside a big rectangle.

1.  243 goes in the very center region which is common to all 3 circles.
2.  537-243=294 were taking History and Dance but not Math.
3.  501-243=258 were taking Math and Dance but not History.
4.  585-243=342 were taking History and Math but not Dance.
5.  1115-258-243-342=272 were taking Math only.
6.  1087-294-243-342=208 were taking History only.
7.  1025-294-243-258=230 were taking Dance only.
8.  2175-230-294-208-258-243-342-272=328 were not taking any of the three.
[These 328 go outside all three circles but inside the rectangle.

Put these 8 results in the 8 regions of a Venn diagram with three overlapping
circles but inside the big rectangle.  Then you'll be ready to answer any
question they can give you.

Edwin


Answer by ikleyn(52781)

.
A group of 2175 students were surveyed about the
courses they were taking at their college with
the following results:
1025 students said they were taking Dance.
1087 students said they were taking History.
1115 students said they were taking Math.
585 students said they were taking History and Math.
501 students said they were taking Math and Dance.
537 students said they were taking History and Dance.
243 students said they were taking all three courses.
~~~~~~~~~~~~~~~~~~~~~~


I read attentively your post, but did not find any question in it.

        Are you sane ?




/////////////////////


Edwin,  there are minimal basic elementary rules of communicating at this forum.

First rule is that a post should be written clearly and correctly,
and the problem should be sensible  (in opposite to be non-sensible).

Second rule is that a post must be complete and contain all necessary data.

Third rule is that a post must contain a question.

Fourth rule is that a post should contain reasonable finite number of questions
(the best one question,  the most - three inter-related questions,  but not 6 or 8 questions).


These rules are simple and necessary,  and  I  ALWAYS  will insist that they be followed.


You,  from your side,  apply 1000-ton efforts to break them and to put me down.


Please stop breaking these rules.   Also,  please stop putting me down.


Next,  why do you think that they want to learn from you about  Venn diagram ?

May be,  in opposite,  they want to learn from  ME  about  Inclusion-Exclusion principle ?


I am the  ONLY  ONE  person at this forum,  who supports some order here,  during the years.

Without me,  it will be  MESS  in two weeks.


Did I say  " in two weeks " ? -   No,  not in two weeks. -   In four days.



Answer by Edwin McCravy(20055)


Stop putting students down, Ikleyn!! They're struggling! Of course this 
student is sane, silly. The main thing was to get the Venn diagram.  That 
may have been all he/she wanted.  It is certainly enough.  There may have 
been several questions but any of them would be trivial once the Venn 
diagram was set up.


Answer by greenestamps(13200)



To tutor @ikleyn....

How absurd -- to claim that if you weren't answering questions on this forum things would be chaos within four days. Such unmitigated arrogance.

And you ABSOLUTELY DESERVE to be put down if you post a response "ARE YOU SANE" because the student neglected to ask a question. Most tutors at this forum who are of better character than you would simply point out to the student that their post contained no question.


Answer by MathTherapy(10552)

A group of 2175 students were surveyed about the 
courses they were taking at their college with 
the following results:

1025 students said they were taking Dance.
1087 students said they were taking History.
1115 students said they were taking Math.
585 students said they were taking History and Math.
501 students said they were taking Math and Dance.
537 students said they were taking History and Dance.
243 students said they were taking all three courses.

It's extremely IRRITATING as well as ANNOYING when one seeks an answer to a mathematics problem and not state
a question or directive in order to get that answer. I usually ignore these posts. However, I don't mind making
an exception and respond to this one.

Here's the Venn
.

If you can read Venn diagrams then you shouldn't have a problem with this one. Nonetheless, I'll 
clarify a few things.

                   M denotes Math
                   D denotes Dance
                   H denotes History

                   Those taking M, only:   272
                   Those taking M&D, only: 258
                   Those taking M&H, only: 342
                   Those taking M,D&H:     243

                   Those taking D, only:   230
                   Those taking D&H, only: 294
                   Those taking H, only:   208
TOTAL (those taking @ least 1 of the 3): 1,847

Those taking NONE of the 3: 2,175 - 1,847 = 328

With all this info. I agree with SIR EDWIN 100%, in that any question can be answered. Then again, and as SIR
EDWIN also pointed out, the Venn diagram may've been all that's needed. If so, that should've been stated. 

SELAH!